# A new generation of effective core potentials: Selected lanthanides and heavy elements II

### v1

@@ III.4 Selected elements
Ru is an element that we have used in calculations before[57]; however, in the past we simply smoothed out the Coulomb cusp present in the MDFSTU ECP. We extended our efforts to optimize the full ccECP that performs comparably to CRENBL, which is the best performing legacy ECP in the atomic spectrum, with CRENBL slightly better in LMAD and ccECP slightly better in overall MAD. The accuracy of the UC atomic spectrum is notably high, making it a challenging benchmark to match. Conversely, its performance in RuO (Fig. 13(b)) shows underbinding at compressed bond lengths, whereas the ccECP performs similarly to CRENBL and MDFSTU. In RuH (Fig. 13(a)), ccECP performs within the chemical accuracy for most bond lengths and matches CRENBL and MDFSTU at the equilibrium bond length.
@@
As mentioned, Cd core is [[Ar]3] and that leaves 20 electrons in the valence space. This makes accurate construction relatively straightforward. The results of the first optimizations outperformed all of the legacy ECPs with the same cores for the atomic spectrum and upon further investigation it also performed very well in the molecular binding tests (Fig. 14). The UC setting slightly outperforms our construction in terms of raw MAD, but the LMAD of ccECP moderately improves upon the UC.
@@ III.5 Selected elements
For the ccECPs we use the same core choice as in our previous work that included selected lanthanide elements [7], [[Kr]4]. This core choice minimizes the number of electrons present in the valence space while at the same time accounts for the effects of tails. This comes to the forefront in hybridizations of molecular or solid state orbitals as well as in formation of large magnetic moments due to partially filled -shell. For these ccECPs we chose to reduce the number of highly ionized states (in the range +3 to +5) in the optimization, as we noticed their inclusion tended to unacceptably bias the overall ccECP quality in bonded environments. We note that CRENBL and LANL2 had different cores for Ce and Eu, hence they were not included in the tests.
@@
The ccECP for lanthanum (La) shows improved performance in atomic spectrum accuracy compared to the other legacy ECPs. We achieve the best values across all metrics, particularly in the LMAD. In LaH (Fig. 15(a)), where the electronic structure is simpler than in LaO (Fig. 15(b)), most ECPs perform within chemical accuracy, out of which the ccECP shows better consistency across the bond lengths tested, especially at equilibrium. For LaO, ccECP reduces the overbinding observed in legacy ECPs for compressed bond lengths, performing within chemical accuracy across all the tested geometries and showing better results than UC at the dissociation limit. During initial optimizations, we encountered challenges in converging states with partially filled orbitals. We have addressed these issues by first optimizing the local channel and subsequently the channel. Once satisfactory performance was achieved for these parameters, we followed the procedure outlined earlier in Sec. II.2.
@@
Cerium (Ce) lies at the beginning of series and has 2 partially filled orbitals, and . This particular configuration posed several challenges in getting the ECP to perform within chemical accuracy in the molecular binding curves and the atomic spectrum simultaneously. During the optimization, we observed the ECPs that were performing extremely well for higher than 3+ ionized states were actually less optimal in molecules. Hence, for better transferability, we included ionized states up to 3+ as well as electron affinity and several excited states. With these considerations we were able to construct a ccECP that outperforms the other legacy ECPs in the atomic spectrum (Figs. 1-3). In the case of molecular binding (Fig. 16), particularly for CeO, the ccECP is the only one that lies within chemical accuracy at equilibrium and extended bond lengths, showing better agreement than the UC method and lying well within the required chemical accuracy for both the molecules being tested. This ensures that the ccECP has better transferability and applicability in bonding and solid state environments without compromising on its performance for low lying states in the atomic system.
@@
As with the previous elements[7], we crafted an initial guess by optimizing the non-local channels independently against , , , and only states corresponding to which channel we were optimizing while a generalized local channel was left fixed. Later we optimized the local and non-locals both against the ground state and its deepest eigenvalues. When doing the final optimizations of the ccECP for europium (Eu) we reduced the number of highly ionized states used in the optimization, focusing on states with 5 ionizations or fewer and several excited states. As seen with previous heavy elements if the ECP is overfit to the atomic spectra, the performance in the molecules drops significantly. We sought to strike a balance with performance in atomic and molecular contexts to maximize transferabilty. As a result, the UC still outperforms the ccECP in the atomic spectrum on some fairly ionized states, but the ccECP is competitive on the low lying states and in the molecules. SBKJC was the only other legacy ECP that shares the same core choice, and while it performs well in EuH, it struggles greatly with EuO, never entering the chemical accuracy regime. EuO has some state crossings that occur in the more compressed region as seen in 17(b). As a result, initially only geometries from Å to Å were converged, but with careful treatment more compressed geometries were also resolved, though there is clearly a discrepancy in which state was converged to. Despite this, judging by the trend in the figure, the discrepancies are small and show no indication that the performance of the ccECP in that region will change meaningfully.
@@ acknowledgments
The data supporting the findings of this study are available in the article and its supplementary material.

### v2

@@ III.4 Selected elements
Ru is an element that we have used in calculations before[75]; however, in the past we simply smoothed out the Coulomb cusp present in the MDFSTU ECP. We extended our efforts to optimize the full ccECP that performs comparably with that of CRENBL, which is the best performing legacy ECP in the atomic spectrum, with CRENBL slightly better in LMAD and ccECP slightly better in overall MAD. The accuracy of the UC atomic spectrum is notably high, making it a challenging benchmark to match. In contrast, its performance in RuO (Fig. 13(b)) shows underbinding at compressed bond lengths, whereas the ccECP performs similarly to CRENBL and MDFSTU. In RuH (Fig. 13(a)), ccECP performs within the chemical accuracy for most bond lengths and matches CRENBL and MDFSTU at the equilibrium bond length.
@@
As mentioned, the Cd core is [[Ar]3] and that leaves 20 electrons in the valence space. This makes accurate construction relatively straightforward. The results of the first optimizations outperformed all legacy ECPs with the same cores for the atomic spectrum, and upon further investigation they also performed very well in the molecular binding tests (Fig. 14). The UC setting slightly outperforms our construction in terms of raw MAD, but the LMAD of ccECP moderately improves upon the UC.
@@ III.5 Selected elements
For the ccECPs we use the same core choice as in our previous work that included selected lanthanide elements [7], [[Kr]4]. This core choice minimizes the number of electrons present in the valence space while at the same time accounting for the effects of tails. This comes to the forefront in hybridizations of molecular or solid-state orbitals as well as in formation of large magnetic moments due to the partially filled -shell. For these ccECPs, we chose to reduce the number of highly ionized states (in the range +3 to +5) in optimization, as we noticed that their inclusion tended to unacceptably bias the overall quality of the ccECP in bonded environments. We note that CRENBL and LANL2 had different cores for Ce and Eu, so they were not included in the tests.
@@
The ccECP for lanthanum (La) shows improved performance in atomic spectrum accuracy compared to the other legacy ECPs. We achieve the best values across all metrics, particularly in the LMAD. In LaH (Fig. 15(a)), where the electronic structure is simpler than in LaO (Fig. 15(b)), most of the ECPs perform within the chemical accuracy, of which the ccECP shows better consistency across the bond lengths tested, especially at equilibrium. For LaO, ccECP reduces the overbinding observed in legacy ECPs for compressed bond lengths, performing within chemical accuracy across all the geometries tested and showing better results than UC at the dissociation limit. During initial optimizations, we encountered challenges in converging states with partially filled orbitals. We have addressed these issues by first optimizing the local channel and then the channel. Once satisfactory performance was achieved for these parameters, we followed the procedure outlined earlier in Sec. II.2.
@@
Cerium (Ce) lies second in the series and has 2 partially filled orbitals, and . This particular configuration posed several challenges in getting the ECP to perform within chemical accuracy in the molecular binding curves and the atomic spectrum simultaneously. During optimization, we observed that ECPs that performed extremely well for higher than 3+ ionized states were actually less optimal in molecules. Hence, for better transferability, we included ionized states up to 3+ as well as electron affinity and several excited states. With these considerations, we were able to construct a ccECP that outperforms the other legacy ECPs in the atomic spectrum (Figs. 1-3). In the case of molecular binding (Fig. 16), particularly for CeO, the ccECP is the only one that lies within chemical accuracy at equilibrium and extended bond lengths, showing better agreement than the UC method and resides well within the required chemical accuracy for both molecules tested. This ensures that the ccECP has better transferability and applicability in bonding and solid-state environments without compromising on its performance for low-lying states in the atomic system.
@@
As with the previous elements[7], we crafted an initial guess by optimizing the non-local channels independently against , , , and only states corresponding to which channel we were optimizing while a generalized local channel was left fixed. Later we optimized the local and non-locals both against the ground state and its deepest eigenvalues. When we preformed the final optimizations of the ccECP for europium (Eu), we reduced the number of highly ionized states used in the optimization, focusing on states with five or fewer ionizations and several excited states. As seen with previous heavy elements, if the ECP is overfit to the atomic spectra, the performance in the molecules drops significantly. We sought to strike a balance with performance in atomic and molecular contexts to maximize transferability. As a result, the UC still outperforms the ccECP in the atomic spectrum on some fairly ionized states, but the ccECP is competitive on the low-lying states and in the molecules. SBKJC was the only other legacy ECP that shares the same core choice, and while it performs well in EuH, it struggles greatly with EuO, never entering the chemical accuracy regime. EuO has some state crossings that occur in the more compressed region as seen in 17(b). As a result, initially only geometries from Å to Å were converged, but with careful treatment, more compressed geometries were also resolved, although there is clearly a discrepancy in which the state was converged. Despite this, judging by the trend in the figure, the discrepancies are small and show no indication that the performance of the ccECP in that region will change meaningfully.
@@ acknowledgments
Author Contributions
Omar Madany: Conceptualization (lead); Investigation (lead); Methodology (equal); Visualization (equal); Writing – original draft (lead); Writing – review & editing (lead). Benjamin Kincaid: Conceptualization (equal); Investigation (equal); Methodology (lead); Visualization (equal); Writing – original draft (supporting); Writing – review & editing (equal); Supervision (lead). Aqsa Shaikh: Conceptualization (supporting); Investigation (equal); Methodology (supporting); Visualization ( supporting); Writing – original draft (supporting); Writing – review & editing (equal). Elizabeth Morningstar: Investigation (supporting). Lubos Mitas: Conceptualization (equal); Investigation (supporting); Methodology (equal); Project administration (lead); Supervision (supporting); Writing – original draft (equal); Writing – review & editing (equal).